# Reassessing Active Learning Adoption in Contemporary NLP: A Community Survey

### v3

@@ Abstract
Supervised learning relies on data annotation which usually is time-consuming and therefore expensive. A longstanding strategy to reduce annotation costs is active learning, an iterative process, in which a human annotates only data instances deemed informative by a model. Research in active learning has made considerable progress, especially with the rise of large language models (LLMs). However, we still know little about how these remarkable advances have translated into real-world applications, or contributed to removing key barriers to active learning adoption. To fill in this gap, we conduct an online survey in the NLP community to collect previously intangible insights on current implementation practices, common obstacles in application, and future prospects in active learning. We also reassess the perceived relevance of data annotation and active learning as fundamental assumptions. Our findings show that data annotation is expected to remain important and active learning to stay relevant while benefiting from LLMs. Consistent with a community survey from over 15 years ago, three key challenges yet persist—setup complexity, uncertain cost reduction, and tooling—for which we propose alleviation strategies. We publish an anonymized version of the dataset. 00footnotetext: Corresponding author: Julia.Romberg@gesis.org [^1] Dataset will be published soon. Equal contribution.
@@ 1 Introduction
Figure 1: Development of AL publication numbers in *CL venues over the past 15 years (total for reference). See Appendix A for methodological details.
@@
Since its beginnings, AL has evolved considerably Xia et al. (2025), not least because of the gains brought about by large language models (LLMs) [^2] We adopt the definition of LLMs from Rogers and Luccioni (2024), which encompasses both encoders and decoders.. We observe notable surges in the number of AL publications in *CL venues in 2018 and 2022 (see Figure 1), likely due to new opportunities enabled by transformer-based language models and generative AI. These works have found considerable improvements in effectiveness and sample-efficiency when integrating LLMs with AL in different scenarios (e.g., Ein-Dor et al., 2020; Margatina et al., 2021; Tonneau et al., 2022; Xiao et al., 2023).
@@
Drawing from the literature alone cannot provide a complete answer to this question, as practical use of AL—in contrast to its technical potential—is often not documented publicly. Therefore, we devise and conduct an extensive community survey following the example of Tomanek and Olsson (2009). While still uncommon in NLP, community surveys have proven effective for obtaining knowledge unavailable in academic literature (e.g., Zhou et al., 2022; Subramonian et al., 2023; Michael et al., 2023; Blaschke et al., 2024). This way, we can not only assess the assumed transformative impact of LLMs in practice, but also take a step back to re-examine whether fundamental assumptions about the need for data-efficient annotation still hold.
@@
(i) We collect a dataset of insights on AL and data annotation needs from the NLP community, academia and industry, through a comprehensive survey with 52 questions. (ii) We evaluate the responses of 144 participants quantitatively and qualitatively, and compare them to the community survey from 2009. (iii) We discuss the results and connect them to literature, describing anticipated future developments in applied AL. (iv) We identify three key problems to adoption that persist in contemporary AL and outline possible solutions.
@@ 2 Survey Methodology
Figure 2: This diagram illustrates (a) the names and sequence of questions groups, (b) the logical flow, including branching points and conditions, and (c) possible survey paths, ordered by the number of questions.
@@ 2.1 Descriptive Research Approach
Following previous community surveys in *CL venues Zhou et al. (2022); Michael et al. (2023); Blaschke et al. (2024), this survey adopts a descriptive research design in order to systematically document experiences within a sample of NLP practitioners. We capture firsthand experiences from academic and industry practitioners relevant to answering the five research questions, including real-world practices, barriers, and further perceptions in the realm of AL adoption. Our primary goal with this approach is to understand how and why practitioners use (or avoid) AL in their specific contexts, instead of making universal claims about AL’s effectiveness. Notably, descriptive research methodology provides valuable documentation, even if it does not generalize to the entire target population. The practitioners’ responses provide empirical evidence that has at most been theoretically discussed, but not systematically documented before. Our findings can be interpreted as valuable documentation of how AL is actually employed in practice, with important implications for future research.
@@ 2.2 Questions and Branching
We use branching logic (Figure 2b) to guide participants through the survey. The main criteria here are having (i) experienced a lack of annotated data, (ii) applied AL in either setting (i.e., simulation or practical), and (iii) used AL in practical settings with human annotators. Questions are automatically skipped if participants are ineligible to answer based on prior responses. For example, participants who have never used AL will not be shown groups III – VI (see Figure 2c for different user groups). Thereby we aim to minimize the required effort for completing the survey, which is more appreciative of the participants’ time, and also has been shown to positively affect the completion rates (Liu and Wronski, 2018). The full survey is provided in Appendix C, including information on each question’s format (i.e., free-form, multiple or single choice).
@@ 2.3 Target Audience and Distribution
We distributed the survey through various channels aiming to reach a broad audience: (1) mailing lists (ACL, ELRA corpora, ELRA SIGUL, Natural Language Processing DC, tada.cool); (2) personalized email to a manually curated list of 601 individuals who co-authored papers on AL at major *CL venues between 2009 and 2024, and further 9 personal contacts; (3) common social media channels (LinkedIn, Bluesky, Twitter/X, Hugging Face Posts); (4) outreach to annotation tool developers and providers, asking them to share the survey with their users. Supplementary details on distribution and implementation are provided in Appendix B.
@@
Figure 3: Exploratory histograms of the participants’ work sectors, field, educational background, work experience, and place of residence (cf. VII.1-5). The x-axis is scaled depending on the smallest and largest percentage value.
@@ 3.1 Participant Statistics
Among 171 persons that navigated to the first question group, 144 completed the survey. [^3] Compared to other surveys Zhou et al. (2022); Subramonian et al. (2023); Michael et al. (2023); Blaschke et al. (2024), 144 full responses place us in the mid range of participation. This corresponds to a completion rate of 84% of those that entered the content part of the survey. Incomplete responses were excluded from analysis.
Figure 3 presents a demographic summary of the 144 participants. Most work in academia, but a substantial portion is employed in industry. Information and communications technology clearly leads the field of work, while the educational background is primarily in computer science and computational linguistics. Overall, the participants are very experienced in NLP, with the majority having between 3 to 10 years of experience. Mailing lists were the most effective referral source (51%, cf. VII.7).
@@ 3.3 Overcoming Data Annotation Challenges
Dissemination of active learning
@@ 3.4 Contemporary Active Learning
Moreover, we ask what contemporary setups of applied AL look like (RQ3). This section evaluates 33 AL projects [^4] There is a near-equal split between academic (16) and industry/government respondents (20, including 3 overlaps). with human annotators that were conducted within the last five years cf. III.5IV.1).
Figure 5: Obstacles especially faced in the application of AL (a) and participants’ general experience with computational methods for a lack of annotated data (b).
Figure 6: Distribution of the reported models and query strategies. Marginal distributions are shown on the side.
@@
We also inquired regarding encountered waiting times and the maximum acceptable waiting times. Contrary to our expectations, participants are willing to wait considerable times up to one week; over 50% even as long as one day (cf. IV.4III.4).
@@
Among those that reported a reduced effectiveness, the main reasons were the lack of suitable annotation tools (24%; notably, 4 had used a tool with AL support) and the overhead of setup (21%), which connects to insufficient knowledge in retrospect (17%). Performance mismatches contributed, such as poor overall performance (21%), a lack of effort reduction (9%), inaccurate efficiency estimation (9%), and mismatching project-specific requirements (6%). In addition, respondents noted sampling bias (15%) and dataset-model dependency (15%). Other options stress concerns about setup time and effectiveness (cf. V.13). Considering potential differences in terms of practical utility “in the real world” and for NLP researchers, we find that academic and industry/governmental respondents align closely on reasons for reduced effectiveness. Appendix D.2 provides a detailed breakdown across sectors for the entire Section 3.4.
@@ 3.5 Anticipated Developments
Answer to RQ4: Practitioners anticipate AL’s future will revolve around pragmatic LLM integration, in particular smaller models. Simultaneously, they call for fundamental improvements in existing query strategies and more accessible tooling.
@@ 4 Related Work
As anticipated in our motivation for conducting this community survey, the existing literature offers only limited answers to the research questions we address. Related work hardly questions the continued need for data annotation (RQ1) and the relevance of AL as a data-efficient method (RQ2). Our paper contributes by providing evidence to support this common assumption. Regarding contemporary AL setups in real-world use (RQ3), reports from actual AL applications or realistic simulations are rare Tonneau et al. (2022). That said, a few studies do address challenges that we identified as persisting in practice, like model-dataset dependency Shelmanov et al. (2021); Tsvigun et al. (2022c); Yüksel et al. (2023); Jelenić et al. (2023) and unstable performance and distribution mismatch Li et al. (2024). The literature provides more extensive coverage of trends (RQ4), complementing the list of potential next steps in applied AL. There is initial research on LLMs as oracles Zhang et al. (2023); Xiao et al. (2023); Liang et al. (2024); Kholodna et al. (2024), AL for few-shot learning or data synthesis Margatina et al. (2023); Bayer and Reuter (2024); Zhang et al. (2023), improving AL on low-resource languages through LLMs Kholodna et al. (2024), cost-efficient LLMs for AL Galimzianova and Sanochkin (2024), LLM-determined curricula for stable AL performance Li et al. (2024), using AL to alleviate LLM hallucinations Xia et al. (2024), handling human label variation Wang and Plank (2023); Van Der Meer et al. (2024) and annotation errors Weber and Plank (2023) with AL.
@@ 5 Discussion
The setup of AL has naturally shifted from maximum-entropy approaches and support-vector machines towards LLMs as the model of choice. At the same time, uncertainty sampling remains the preferred query strategy. Although surprising at first glance, given the growing range of more sophisticated alternatives, this is in accordance with the literature on strong performance of uncertainty sampling Shen et al. (2017); Margatina et al. (2022); Schröder et al. (2022). In terms of stopping criteria, practical constraints such as time and budget still largely outweigh information-theoretical grounds.
@@
While LLMs offer options beyond serving as backbone models, these have not yet been integrated into the AL workflows reported by our participants. Approaches like prompt-based strategies Li et al. (2024) and in-context learning Margatina et al. (2023) could make AL more accessible for less-expert users, and using AL for few-shot example selection Bayer and Reuter (2024) may further reduce annotation effort. Trends such as LLMs-as-annotators, however, require stronger theoretical foundations first before they can provide reliable support in practice.
Despite progress on single issues, such as model–dataset dependency Tsvigun et al. (2022c); Jelenić et al. (2023) and performance stability Li et al. (2024), research overall did not manage to sufficiently address the three overarching challenges to practical adoption that the community has raised. These are: (1) Excessive complexity of the AL setup: Respondents who never had applied AL most often cite a lack of methodological expertise and implementation overhead as obstructing factors. AL users share the concerns regarding methodological choices. The ongoing introduction of new algorithms (e.g., query strategies) and, thus, possible setup combinations continually increases complexity. (2) Risk of starting an AL project: Respondents noted that the cost reduction through AL cannot be reliably estimated upfront. Its dependence on numerous variables makes a priori chosen setups prone to unsatisfactory outcomes Lowell et al. (2019), possibly worse than random sampling Ghose and Nguyen (2024). (3) Perceived insufficiency of tooling: Despite considerable activity in AL tooling, it is still perceived as insufficient. We suspect this is partly due to the large problem space, making it difficult for tools to serve all tasks.
As a starting point for addressing these challenges, we propose to: (1) prioritize the reduction of existing complexity instead of devising new algorithms, e.g., by refuting the effectiveness of established strategies; (2) develop and continually research guidelines or heuristics for choosing AL components based on the setup’s variables, e.g., with a decision tree that specifies based on task and number of classes. More insights from practical use cases would also be immensely helpful to provide valuable points of comparison; and (3) enhance the visibility of existing tools, and provide user feedback that can be used by tool developers to refine their interfaces to ease the application of AL Shnarch et al. (2022b); Jukić et al. (2023), especially for non-experts Ras et al. (2022).
@@ 6 Conclusions
It is difficult to determine when participants formed their opinions on AL, as assessments are commonly based on past experiences that may not align with current advancements. Similarly, we cannot ensure that recent improvements in AL were acknowledged or influenced their opinion. As a result, the collected responses may have a certain temporal bias, potentially pointing to issues that have since been improved upon or even resolved. However, our analysis of recent projects confirms that the three key challenges (setup complexity, cost reduction estimation, and tooling) persist, to a certain extent.
@@
To comply with data protection laws, we took the necessary steps to ensure compliance with the EU General Data Protection Regulation. Respondents granted their consent to the storage and analysis of their provided data, and the distribution of an anonymized version thereof, before starting the survey. To allow the NLP community for extended analysis of the collected information and to uphold scientific transparency, we will publish the resulting dataset in a form that does not allow to infer conclusions about individuals.

### v4

@@ Abstract
Supervised learning relies on data annotation which usually is time-consuming and therefore expensive. A longstanding strategy to reduce annotation costs is active learning, an iterative process, in which a human annotates only data instances deemed informative by a model. Research in active learning has made considerable progress, especially with the rise of large language models (LLMs). However, we still know little about how these remarkable advances have translated into real-world applications, or contributed to removing key barriers to active learning adoption. To fill in this gap, we conduct an online survey in the NLP community to collect previously intangible insights on current implementation practices, common obstacles in application, and future prospects in active learning. We also reassess the perceived relevance of data annotation and active learning as fundamental assumptions. Our findings show that data annotation is expected to remain important and active learning to stay relevant while benefiting from LLMs. Consistent with a community survey from over 15 years ago, three key challenges yet persist—setup complexity, uncertain cost reduction, and tooling—for which we propose alleviation strategies. We publish an anonymized version of the dataset.00footnotetext: Equal contribution.00footnotetext: Corresponding author: Julia.Romberg@gesis.org [^1] Dataset is available from: https://doi.org/10.7802/2990.
@@ 1 Introduction
Figure 1: Development of AL publication numbers in *CL venues over the past 15 years (total publication numbers for reference). The methodological details are provided in Appendix A. Note that the numbers for 2025 must be interpreted with caution, as not all data were available in the ACL Anthology at the time of analysis.
@@
Since its beginnings, AL has evolved considerably Xia et al. (2025), not least because of the gains brought about by large language models (LLMs [^2] We adopt the definition of LLMs from Rogers and Luccioni (2024), which encompasses both encoders and decoders.). We observe notable surges in the number of AL publications in *CL venues in 2018 and 2022 (see Figure 1), likely due to new opportunities enabled by transformer-based language models and generative AI. These works have found considerable improvements in effectiveness and sample efficiency when integrating LLMs with AL in different scenarios (e.g., Ein-Dor et al., 2020; Margatina et al., 2021; Tonneau et al., 2022; Xiao et al., 2023).
@@
Drawing from the literature alone cannot provide a complete answer to this question, as practical use of AL—in contrast to experimental evaluations—is often not documented publicly. Therefore, we devise and conduct an extensive community survey following the example of Tomanek and Olsson (2009). While still uncommon in NLP, community surveys have proven effective for obtaining knowledge unavailable in academic literature (e.g., Zhou et al., 2022; Subramonian et al., 2023; Michael et al., 2023; Blaschke et al., 2024). This way, we can not only assess the assumed transformative impact of LLMs, but also take a step back to re-examine whether fundamental assumptions about the need for data-efficient annotation still hold.
@@
(1) We collect a dataset of insights on AL and data annotation needs from the NLP community, academia and industry, through a comprehensive survey with 52 questions. (2) We evaluate the responses of 144 participants quantitatively and qualitatively, and compare them to the community survey from 2009. (3) We discuss the results and connect them to literature, describing anticipated future developments in applied AL. (4) We identify three key problems to adoption that persist in contemporary AL and outline possible solutions.
@@ 2 Survey Methodology
Figure 2: This diagram illustrates (a) the names and sequence of question groups in the survey, (b) the logical flow between these question groups, including branching points and conditions, and (c) the possible paths that respondents take through the survey, ordered by the number of questions in descending order.
@@ 2.1 Descriptive Research Approach
Following previous community surveys in *CL venues Zhou et al. (2022); Michael et al. (2023); Blaschke et al. (2024), this survey adopts a descriptive research design in order to systematically document experiences within a sample of NLP practitioners. We capture firsthand experiences from academic and industry practitioners relevant to answering the five research questions, including real-world practices, barriers, and further perceptions of AL adoption. Our primary goal with this approach is to understand how and why practitioners use (or avoid) AL in their specific contexts, instead of making universal claims about AL’s effectiveness. Notably, descriptive research methodology provides valuable documentation, even if it does not generalize to the entire target population. The practitioners’ responses provide empirical evidence that has at most been theoretically discussed, but not systematically documented before. Our findings can be interpreted as valuable documentation of how AL is actually employed in practice, with important implications for future research.
@@ 2.2 Questions and Branching
We use branching logic (Figure 2b) to guide participants through the survey. The main criteria here are having (1) experienced a lack of annotated data, (2) applied AL in either setting (i.e., simulation or practical), and (3) used AL in practical settings with human annotators. Questions are automatically skipped if participants are ineligible to answer based on prior responses. For example, participants who have never used AL will not be shown groups III–VI (see Figure 2c for different user groups). Thereby we aim to minimize the required effort for completing the survey, which is more appreciative of the participants’ time, and also has been shown to positively affect the completion rates (Liu and Wronski, 2018). The full survey is provided in Appendix C, including information on each question’s format (i.e., free-form, multiple or single choice).
@@ 2.3 Target Audience and Distribution
We distributed the survey through various channels aiming to reach a broad audience: (1) mailing lists (ACL, ELRA corpora, ELRA SIGUL, Natural Language Processing DC, tada.cool); (2) personalized email to a manually curated list of 601 individuals who co-authored papers on AL at major *CL venues between 2009 and 2024, and 9 additional personal contacts; (3) common social media channels (LinkedIn, Bluesky, Twitter/X, Hugging Face Posts); (4) outreach to annotation tool developers and providers, asking them to share the survey with their users. Supplementary details on distribution and implementation are provided in Appendix B.
@@
Figure 3: Exploratory histograms of participants’ work sectors, the field they are working in, educational background, work experience in machine learning/NLP, and primary place of residence (cf. VII.1-5). The y-axis of each histogram represents the response options for the survey question, while the x-axis shows the proportion of participants who chose each response option.
@@ 3.1 Participant Statistics
Among 171 persons that navigated to the first question group, 144 completed the survey. [^3] Compared to other surveys Zhou et al. (2022); Subramonian et al. (2023); Michael et al. (2023); Blaschke et al. (2024), 144 full responses place us in the mid-range of participation. This corresponds to a completion rate of 84% of those that entered the content part of the survey. Incomplete responses were excluded from analysis.
Figure 3 presents a demographic summary of the 144 participants. Most work in academia, but a substantial portion is employed in industry. Information and communications technology clearly leads the field of work, while the educational background is primarily in computer science and computational linguistics. Overall, the participants are very experienced in NLP, with the majority having between 3 to 10 years of experience. For recruiting participants, mailing lists were the most effective referral source (51%, cf. VII.7).
@@ 3.3 Overcoming Data Annotation Challenges
Figure 5: Obstacles especially faced in the application of AL (a) and participants’ general experience with computational methods for a lack of annotated data (b). The y-axis represents the response options for the survey question, while the x-axis shows the proportion of participants who chose each response option.
Adoption of active learning
@@ 3.4 Contemporary Active Learning
Moreover, we ask what contemporary setups of applied AL look like (RQ3). This section evaluates 33 AL projects [^4] There is a near-equal split between academic (16) and industry/government respondents (20, including 3 overlaps). with human annotators that were conducted within the last five years (from 2020 onwards; cf. III.5IV.1).
Figure 6: Distribution of the reported combinations of model and query strategy.
Figure 7: Waiting times between two annotation cycles encountered in practical projects (cf. IV.4) versus the maximum time respondents would accept (cf. III.4).
@@
We also inquired regarding encountered waiting times and the maximum acceptable waiting times between annotation cycles, as illustrated in Figure 7. We observe a broad range reported across projects, from under one minute to up to one week. Contrary to our expectations, participants are willing to wait considerable times up to one week, over 50% even as long as one day, suggesting that significant sampling times are neither unexpected nor problematic for them.
@@
Among those that reported a reduced effectiveness, the main reasons were the lack of suitable annotation tools (24%; notably, 4 had used a tool with AL support) and the overhead of setup (21%), which connects to insufficient knowledge in retrospect (17%). Performance mismatches contributed, such as poor overall performance (21%), a lack of effort reduction (9%), inaccurate efficiency estimation (9%), and mismatching project-specific requirements (6%). In addition, respondents noted sampling bias (15%) and dataset-model dependency (15%). The other options stress concerns about setup time and effectiveness (cf. V.13). Considering potential differences in terms of practical utility “in the real world” and for NLP researchers, we find that academic and industry/governmental respondents align closely on reasons for reduced effectiveness. Appendix D.2 provides a detailed breakdown across sectors for the entire Section 3.4.
@@ 3.5 Anticipated Developments
Answer to RQ4: Practitioners anticipate that the future of AL will revolve around increased and pragmatic LLM integration. Simultaneously, they call for fundamental improvements in existing query strategies and more accessible tooling.
@@ 4 Related Work
As anticipated in our motivation for conducting this community survey, the existing literature offers only limited answers to the research questions we address. Related work hardly questions the continued need for data annotation (RQ1) and the relevance of AL as a data-efficient method (RQ2). Our paper contributes by providing evidence to support this common assumption. Regarding contemporary AL setups in real-world use (RQ3), reports from actual AL applications or realistic simulations are rare Tonneau et al. (2022). That said, a few studies do address challenges that we identified as persisting in practice, such as model-dataset dependency Shelmanov et al. (2021); Tsvigun et al. (2022c); Yüksel et al. (2023); Jelenić et al. (2023) and unstable performance and distribution mismatch Li et al. (2024). The literature provides more extensive coverage of trends (RQ4), complementing the list of potential next steps in applied AL. There is initial research on LLMs as annotators Zhang et al. (2023); Xiao et al. (2023); Liang et al. (2024); Kholodna et al. (2024), AL for few-shot learning or data synthesis Margatina et al. (2023); Bayer and Reuter (2024); Zhang et al. (2023), improving AL on low-resource languages through LLMs Kholodna et al. (2024), cost-efficient LLMs for AL Galimzianova and Sanochkin (2024), LLM-determined curricula for stable AL performance Li et al. (2024), using AL to alleviate LLM hallucinations Xia et al. (2024), handling human label variation Wang and Plank (2023); Van Der Meer et al. (2024) and annotation errors Weber and Plank (2023) with AL.
@@ 5 Discussion
The setup of AL has naturally shifted from maximum entropy approaches and support vector machines towards LLMs as the model of choice. Our survey indicates the popularity of smaller LLMs (in this study defined as models with less than 1B parameters). We hypothesize that the close or even on par performance these models can achieve with fine-tuning on many traditional supervised tasks (specifically encoders), combined with their lower hardware requirements, render larger models (i.e., LLMs with 1B parameters or more) computationally inefficient for practical application in AL. Larger LLMs may become the preferred choice in cases where a potentially minor performance increase is mission critical, or when no suitable smaller model exists (e.g., for languages where no strong encoder models are available). At the same time, uncertainty sampling remains the preferred query strategy. Although surprising at first glance, given the growing range of more sophisticated alternatives, this is in accordance with the literature on strong performance of uncertainty sampling Shen et al. (2017); Margatina et al. (2022); Schröder et al. (2022). In terms of stopping criteria, practical constraints such as time and budget still largely outweigh information-theoretical grounds.
@@
While LLMs offer options beyond serving as backbone models, these have not yet been integrated into the AL workflows reported by our participants. Approaches like prompt-based strategies Li et al. (2024) and in-context learning Margatina et al. (2023) could make AL more accessible for less-expert users, and using AL for few-shot example selection Bayer and Reuter (2024) may further reduce annotation effort. Trends such as LLMs-as-annotators, however, require stronger theoretical foundations first before they can provide reliable support in practice Baumann et al. (2025); Tan et al. (2024).
Despite progress on single issues, such as model-dataset dependency Tsvigun et al. (2022c); Jelenić et al. (2023) and performance stability Li et al. (2024), research overall did not manage to sufficiently address the three overarching challenges to practical adoption that the community has raised:
  1. (1)
    Excessive complexity of the AL setup: Respondents who never had applied AL most often cite a lack of methodological expertise and implementation overhead as obstructing factors. AL users share the concerns regarding methodological choices. The ongoing introduction of new algorithms (e.g., query strategies) and, thus, possible setup combinations continually increases complexity.
  2. (2)
    Risk of starting an AL project: Respondents noted that the cost reduction through AL cannot be reliably estimated upfront. Its dependence on numerous variables makes a priori chosen setups prone to unsatisfactory outcomes Lowell et al. (2019), possibly worse than random sampling Ghose and Nguyen (2024).
  3. (3)
    Perceived insufficiency of tooling: Despite considerable activity in the development of AL annotation tools, their progress is still perceived as insufficient. We suspect that this is partly due to the large problem space, making it difficult for tools to serve all tasks.
As a starting point for addressing these challenges, we propose to:
  1. (1)
    prioritize the reduction of existing complexity instead of devising new algorithms, e.g., by refuting the effectiveness of established strategies;
  2. (2)
    develop and continually research guidelines or heuristics for choosing AL components based on the setup’s variables, e.g., with a decision tree that specifies based on task and number of classes. More insights from practical use cases would also be immensely helpful to provide valuable points of comparison; and
  3. (3)
    enhance the visibility of existing tools, and provide user feedback that can be used by tool developers to refine their interfaces to ease the application of AL Shnarch et al. (2022b); Jukić et al. (2023), especially for non-expert users Ras et al. (2022).
@@ 6 Conclusions
It is difficult to determine when participants formed their opinions on AL, as assessments are commonly based on past experiences that may not align with current advancements. Similarly, we cannot ensure that recent improvements in AL were acknowledged or influenced their opinion. As a result, the collected responses may be subject to a temporal bias, potentially pointing to issues that have since been improved upon or even resolved. However, our analysis of recent projects confirms that the three key challenges (setup complexity, cost reduction estimation, and tooling) persist, to a certain extent.
@@
To comply with data protection laws, we took the necessary steps to ensure compliance with the EU General Data Protection Regulation. Respondents granted their consent to the storage and analysis of their provided data, and the distribution of an anonymized version thereof, before starting the survey. To allow the NLP community to conduct extended analyses on the collected information and to uphold scientific transparency, we will publish the resulting dataset in a form that does not allow to infer conclusions about individuals.